It would be greatly appreciated if you would be able to help me with:
The sum of squares of two positive consecutive odd integers is 74. Find the integers.
I've made an attempt at:
(x^2)+(x+2)74
Thanks for your help!

Answer by rothauserc(4718)

let x be the first odd integer, then x+2 is the next consecutive odd integer
(x^2) + (x+2)^2 = 74
x^2 + x^2 + 4x +4 = 74
2x^2 + 4x - 70 = 0
(2x-10)*(x+7) = 0
we have two solution sets
x = 5 or -7
our solution sets are
5, 7 or -7, -5
check the answers
5^2 + 7^2 = 25 + 49 = 74
(-7)^2 + (-5)^2 = 49 + 25 = 74
our answers check :-)